ON THE OUTCOME OF THE REACTIONS BETWEEN HYDROCARBON RADICALS AND O_{2} IN HELIUM NANODROPLETS

\underline{A. M. MORRISON} and G. E. DOUBERLY, Department of Chemistry, University of Georgia, Athens, Georgia 30602-2556.

Helium nanodroplet isolation and infrared laser spectroscopy are used to investigate the CH_{3} + O_{2} and C_{3}H_{3} (propargyl) + O_{2} reactions. The hydrocarbon radicals are generated in an effusive pyrolysis source located upstream from a differentially pumped O_{2} gas pick-up cell. In this experimental configuration, the reaction occurs between sequentially picked-up and presumably cold fragments. The CH_{3} + O_{2} reaction leads barrierlessly to the methyl-peroxy radical, and despite having to dissipate an energy of approximately 30 kcal/mol, the infrared spectra reveal a large abundance of droplets containing the cold CH_{3}O_{2} radical. Theoretical studies have predicted an approximately 2-4 kcal/mol barrier in the entrance channel of the C_{3}H_{3} + O_{2} reaction. Therefore, we initially expected to see a weakly bound entrance channel C_{3}H_{3}–O_{2} van-der-Waals complex, given the rapid cooling provided by the dissipative helium environment. However, only the trans-acetylenic isomer of the propargyl-peroxy radical is observed. The dipole moment of this species is measured with infrared laser Stark spectroscopy.